OpenStudy (anonymous):

Simplify, assuming that the factors are approximations. Give answers in scientific notation with the same number of significant digits as in t he least accurate factor. (8 x 10^6) x (2 x 10^-2) / 4 x 10^2

pooja195 (pooja195):

8*10^6 * ( 2* (1/10^2)) --------------------- = 2*10^4 * 2/10^2 = 4*10^2 4*10^2

pooja195 (pooja195):

@RH do u understand?????????

OpenStudy (anonymous):

YES! Thank you very much :)